MICQRC(5) MICQRC(5) NAME ˜ /.micq/micqrc−Theconfigurationfileformicq DESCRIPTION Thismanualpagedescribestheconfigurationfilesformicq.Itconsistsofseveralsectionsthatare startedby[section],wheresectionisthenameofthesection.Fileordirectorynamesgivenare takenrelativetothebasedirectory,whichisusually˜ /.micq/. CONNECTIONS AConnectionsectiondescribesaconnectionwithinICQ.Typically,therearetwoofthem,that arerelatedtoeachother:aconnectiontotheICQserver,andaconnectiontolistentopeerto peercommunication.Youmayalsohavearemotecontrolconnection:thatmeansaFIFOis opened,andlinesreadfromitareinterpretedinthesamewayasinputlines. type<type>[auto] Settheconnectiontypetotype,whichcanbeicq5foraconnectiontoanICQv5server, icq8foraconnectiontoanICQv8server,peerforaconnectionlisteningtopeertopeer communication,orremoteforaremotecontrolFIFO.Ifautoisgiven,createthisconnectiononstartup.Youpresumablewanttodothis,unlessyouhaveseveralUINs,then setthisontheconnectioncorrespondingtothisUINyouareusing. version<nr> Settheversionoftheprotocolusedforthisconnectiontonr.Pleasenotethatfora serverconnections,versions5,7and8aresupported,andforpeertopeerconnections6, 7and8.RemotecontrolFIFOsdon´ thaveaversion. server<server> Settheservertoconnecttotoserver.Thisdefaultstologin.icq.com.Unusedforpeerto peerconnections.Forremotecontrolfifos,thisisthepathoftheFIFOtobecreated. port<port> Settheporttoconnectontheservertoport,orthefirstporttotrytolistenonforpeer topeerconnectionstoport. uin<uin> SettheUINtologinwithtouin.UnusedforpeertopeerconnectionsandremotecontrolFIFOs. password[<password>] Setthepasswordtologinwithtopassword.Unusedforpeertopeerconnections.If omitted,askforpasswordonstartup.Note:BeOSusersmaynotomitthepassword. status<nr> Setthestatustobeonloginintonr.Statuscanbe: 0online 1away 5notavailable 17occupied 19donotdisturb 32freeforchat 256invisible Forpeertopeerconnections,thisisusedforthedirectconnectionsetting: 1We´ refirewalledandcan´ treceiveanyconnection.Don´ teventry. 2We´ refirewalled,butuseSOCKSandcanreceiveconnections. mICQ 1 MICQRC(5) MICQRC(5) 4We´ renotfirewalledatall.Thisisthedefault. RemotecontrolFIFOsdonotusethissetting. GENERAL TheGeneralsectionincludestheoptionsforSOCKS5,thatmaybeeventuallymovedtotheir ownsection,generaloptions,andsomestringstooverride. s5use<number> Setnumberto0todisableand1toenable. s5host<host> Connecttosocks5proxyonhosthost. s5port<port> Connecttosocks5proxylisteningonportport. s5auth<number> Setnumberto0toomitauthenticationand1todoauthentication. s5name<name> Identifyasusernametosocks5proxy. s5pass<password> Authenticatewithpasswordpasswordtosocks5proxy. Afewbook-keepingoptionsareusuallyplacedatthebeginningofthemicqrcfile: encodingfile<encoding> Specifytheencodingwhichisusedforstringsinthisfileasencoding.mICQwillalways writethisfileinUTF-8. encodinglocal<encoding> Specifytheuser´ sterminal´ sencodingasencoding.Thisoverridestheencodingderived fromenvironmentvariables. Note:Usingthiswillcausethelocaletobeflaggedasbroken,thuscausingcharacters outsidetheiso-8859-1rangetobeprintedininputhandlingashexunicodecodepointsas theirscreencellwidthcannotbedeterminedwithwcwidth()anymore.Thus,alwaysuse theLCCTYPElocalesettingtooverridetheusedcharacterencoding. encodingremote<encoding> Settheencodingtouseformetadataontheserver,includingthepassword,toencoding. Thisencodingisalsousedforcommunicatingwithothercontactswhoseclientdoesn´ t understandutf8messagesunlessaspecificencodingissetforthiscontact.Deprecatedas ofmICQ0.5;usetheencodingglobaloptioninstead. format<nr> Settheformatthisconfigfileistonr.Format2iscurrent,butformat1isalsoaccepted. Nowtheotheroptions: verbose<level> Settheverbositylevelonstartuptolevel. autoaway<time> Setthetimetosetstatustoawayaftertotimeseconds.Setto0todisable. Screenwidth<width> Setthedefaultscreenwithtowidth.Ifunsetor0,80willbeused.Thisisrarelyneeded asusuallythecurrentscreenwidthcanbededucedatruntime. setdelbson|off Ifsettoon,letthedeletecharacteractasabackspace.Whenoff,letitactasadelete character. mICQ 2 MICQRC(5) MICQRC(5) setfunnyon|off Disableorenable"funny"messages. setcoloron|off Disableorenabletheuseofcolors. sethermiton|off Ifsettoon,ignoreallmessagesfromeveryonenotonyourcontactlist.Deprecatedasof mICQ0.5,settheglobaloptionignoretoonforthesameeffect. setlogon|off Disableorenablelogging.DeprecatedasofmICQ0.5,settheglobaloptionlogmessto onforthesameeffect. setlogonoffon|off Whenlogging,disableorenabletheloggingofonline/offlinestatuschanges.Deprecated asofmICQ0.5,settheglobaloptionslogmessandlogchangetoonforthesameeffect. setautoon|off Disableorenableautomaticresponses(replymessages). setuinprompton|off Disableorenabletheuseofthenickyoulastreceivedamessagefromintheprompt. setautosaveon|off Disableorenabletheautomaticsavingofthemicqrc. setautofingeron|off DisableorenabletheautomaticfingeringofnewUINs. setlinebreak<type> Configurethelinebreaktypeforincomingmessagestotype,whichcanbesimpletojust wraplinesasusual,breaktohavealinebreakbeforeeachmessage,indenttohavea linebreakandindentthemessageandsmarttohavealinebreakonlyifthemessage doesn´ tfitonthecurrentline. settabssimple|cycle|cycleall DeprecatedasofmICQ0.5,asitusesamuchimprovedtabhandlingnowwhichhasallof thepreviousfeatures. setsilent<type> Suppresssomeoutput,namelystatuschangesforonandstatuschanges,loginsand logoutsforcomplete.DeprecatedasofmICQ0.5,useglobaloptions showchangeandshowonofftoonforthesameeffect. options<options> Setglobaloptions.Seetheoptcommandfordetails. chat<nr> Settherandomchatgrouptonr.Use−1todisable,and49formICQ(whichisthe default). autoupdate<level> Currentlevelofautomaticconfigurationupdatesdone.Willbeincrementedtoabigger valueeachtimeconfigurationisupdatedtonewdefaultvalues.Donotmakethemistake tosetto0todisableasthatwillhavetheeffectofallupdatesbeingre-done. Atlast,somestringscanbedefined: colorscheme<nr> Selectthecolorschemenumbernr. color<use><color> Selectcolorcolorforuse.usecanbeanyofnone,server,client,message,contact, sent,ack,error,debugorincoming,whilecolorcanbeanyoneofblack,red,green, mICQ 3 MICQRC(5) MICQRC(5) yellow,blue,magenta,cyan,white,none,orboldoracombinationofthose(bold, however,mustbelasttotakeeffect),oranyverbatimstringtomaketheuser´ sterminal selectthedesiredcolor. logplace<file>|<dir> Setthefiletologintotofile,orthedirectorytologintotodir.Pleasenotethatapath isassumedtobeadirectoryifithasatrailing/. soundon|beep|off|event Specifywhathappensifabeepistobegenerated.onorbeepwillsimplybeep,offwill donoting,whileeventwillcallthescriptforevents. event<script> Setthescripttoexecuteforeventstoscript.Itiscalledwiththefollowingarguments: 1.TheIMtype,currentlyonlyicq. 2.TheUINofthecontactthiseventrelatesto,or0. 3.Thenickofthecontactthiseventstorelatesto,ortheemptystring. 4.Thestringglobal. 5.Thetypeofthisevent,whichcanbemsg,on,off,beeporstatus,whereonandoff areforoncomingandoffgoingcontacts.Thislistmaynotbeexhaustive. 6.Formessages,themessagetype,foroffgoingcontacts,thepreviousstatus,andfor oncomingcontactsorstatuschanges,thenewstatus,otherwise0. 7.Thetextofthemessage. 8.Thecontact´ suseragent. Notethatforsecurityreasons,singlequotesmaybereplacedbydoublequotes,andthe messagetextmaybetruncated. auto<status><string> Settheautomaticreplyinstatusstatustostring.Thisoptionmayberepeatedforall possiblevaluesaway,na,dnd,occ,inv,andffcforstatus. STRINGS TheStringssectioncontainscommandrenames. alter<old><new> Renamecommandoldtonew.Notetheoldnamemaystillbeused,unlessitconflicts withsomenewname.Forpossiblecommandnames,seemicq(7).Thisoptionmaybe repeatedasdesired. Note:thisoptionisobsolete,usealiasinstead. alias[auto[expand]]<alias><expansion> Defineanaliasnamedaliaswhichissubstitutedwithexpansion.Iftheautoorautoexpandkeywordisgiven,theexpansionalsohappenswhenthespacekeyorthereturnkey ispresseddirectlyafterthealias.Ifthestring%sispresentinexpansion,eachoccurence isreplacedwiththegivenargumentswhenthealiasisinvoked,otherwisetheywillbe appended.Ifthestring%rispresentinexpansion,eachoccurenceisreplacedbythe nicknameor,ifthenickisnotavailable,theUIN,ofthecontactthatsentthelastmessagetoyou,ortheemptystringifthereisnone.Ifthestring%aispresentinexpansion, eachoccurenceisreplacedbythenicknameor,ifthenickisnotavailable,theUIN,of thecontactthatyousentthelastmessageto,ortheemptystringifthereisnone. Newfor0.4.10.Auto-expandingaliasesnewfor0.5.0.4. GROUP TheGroupsectioncontactsacontactgroupandmayberepeatedasrequired.Notethatforformat2,itissavedintheBASE/statusfile,notinBASE/micqrc.Itmayhavethefollowingcommands: server<type><uin> Settheserverconnectionthiscontactgroupbelongstototheoneoftypetype,whichcurrentlycanbeicqv8forICQserverconnectionsversion8oricqv5forICQserver mICQ 4 MICQRC(5) MICQRC(5) connectionsversion5,andforUINuin.Ifthiscommandisomitted,thefirstactive serverconnectionisassumed. label<label> Setthelabelofthiscontactgrouptolabel.Ifitiscontacts-<type>-<uin>,thenitisthe contactlistforthisserverconnection.Notethatforformat2oftheBASE/statusfile,this isobsolete,asthecontactsdistributedtotheserverconnectionsaccordingtotheserver statementintheirContactssection. id<id> Settheidofthiscontactgrouptoid.0meansnoiddefined.Thiswillbesetwhen downloadingoruploadingacontactlist. options<options>... Defineoptionsforthiscontactgroup.Seetheoptcommandfordetails. entry<id><uin> AddcontactwithUINuinasididtothiscontactgroup.Notethatidisobsoleteasof mICQ0.5andalwayssetto0.TheidissetintheContactssectioninstead. CONTACTS(obsoleteformat) TheContactssectioncontainsthe(global)contactlist.Itisobsoleteinthisformat. [*][˜ ][ˆ ]uinnick MakeuserwithUINuinknownunderthenicknick.If*isgiven,theusermayseeyou whileyou´ reinvisible.If˜ isgiven,lethimseeyouasalwaysoffline.Ifˆ isgiven,ignore thisuser.Ifauinoccursmultipletimes,allexceptthefirstaretreatedasaliases. CONTACTS TheContactssectioncontainsacontactlistforoneserverconnection.Itisrepeatedforeach serverconnection.NotethatitissavedintheBASE/statusfile,notinBASE/micqrc.Itmay havethefollowingcommands: server<type><uin> Definestheserverconnectionthisisthecontactlistfor,withthesamesyntaxasinthe Groupsection. entry<id><uin><nick>... AddsuserwiththeUINuinandididtothiscontactlist,withnickandarbitrarymany aliasesnick. options<options>... Defineoptionsforthepreviouscontact.Seetheoptcommandfordetails. SEEALSO micq(1),micq(7) AUTHOR ThismanpagewaswrittenbyJamesMorrison<ja2morrison@student.math.uwaterloo.ca>.Itwas rewrittentomatchnewconfigfilesyntaxbyRu¨ digerKuhlmann<micq@ruediger-kuhlmann.de>. mICQ 5